OpenStudy (anonymous):

Let v1 = (-6,4) and v2 = (-3, 6). how to find the scalar projection of v1 onto v2?

OpenStudy (anonymous):

\[proj_{V _{2}}V _{1}=\frac{ V _{1}.V _{2} }{ V _{2} . V _{2}}V _{2}\]

OpenStudy (anonymous):

So that is the dot product of v1 and v2 divided by v2 dotted with v2... all times v2

OpenStudy (anonymous):

Do you know dot products?

OpenStudy (anonymous):

not exactly but i can try with the information you gave me.

OpenStudy (nurali):

The scalar (or component) of v1 onto v2 is comp_v1(v2) = v1∙v2/||v1|| = (-6(-3) + (4)(6))/√((-6)²+(4)²) = 5.82

OpenStudy (anonymous):

\[(a,b).(c,d)=ac+bd\] you multiply corresponding entries and then add up those products.

OpenStudy (anonymous):

thanks. (: mind helping with one last question related to this problem?

OpenStudy (anonymous):

not scalar projection but Just the projection of v1 onto v2

OpenStudy (nurali):

The projection of v1 onto v2 is simply the scalar component times a unit vector in the direction of v1. proj_v1(v2) = v1∙v2/||v1||² v1 = (14/15)(-6,4) = (-28/5, 56/15)

OpenStudy (anonymous):

Actually I should have read THIS problem better. The SCALAR projection is what I said but without the multiplication by v2. What I described is projection (aka vector projection)

OpenStudy (anonymous):

What Nurali put is the scalar projection.

OpenStudy (anonymous):

okay, I see. thanks for all the help(:

OpenStudy (nurali):

The projection of v1 onto v2 is simply the scalar component times a unit vector in the direction of v1. proj_v1(v2) = v1∙v2/||v1||² v1 = (21/26)(-6,4) = (-63/13, 42/13)

OpenStudy (nurali):

My Pleasure.